OpenStudy (anonymous):

The ratio of the heights of two similar rectangular prisms is 2:3. What is the ratio of their lateral areas? 8:27 2:5 4:9 1:0.33

OpenStudy (anonymous):

its 4:9

OpenStudy (anonymous):

A rectangular prism is truncated removing 10% of its height. The original height of the prism was 100 meters and it has a length of 4 meters and a width of 20 meters. What is the volume of the truncated prism? 800 m3 4,200 m3 3,800 m3 7,200 m3

OpenStudy (anonymous):

vol = l*w*h vol = 4*20*100 vol = 8000 m^3

OpenStudy (anonymous):

as we are reducing 10% of its height so h= 90 so vol = 7,200m^3

OpenStudy (anonymous):

answer is 7200m^3

OpenStudy (anonymous):

hope you got it :)

OpenStudy (anonymous):

What is the similarity ratio of two cubes with volumes of 125 cubic meters and 216 cubic meters? 125:216 25:36 1:3 5:6

OpenStudy (anonymous):

answer is 5:6

OpenStudy (anonymous):

last one :) Part 1: Create and provide the dimensions for two similar figures of your choosing. Part 2: What is the similarity ratio of these figures along with the ratio of their surface area and volume? Part 3: Show your work, either using the actual volumes or using the formula, that the volume ratio is true.

OpenStudy (anonymous):

need time for this one

OpenStudy (anonymous):

alright:)

OpenStudy (anonymous):

first part: let, fig 1 fig2 L=2 cm L=4cm W=4 cm W=8cm H=5 cm H=10cm part 2: The first fig is half the size of the second fig so their similarity ratio would be 1:2 Surface Area of a Rectangular Prism = 2lh + 2hw + 2lw fig1 surface area = 75cm^2 fig2 surface area = 304cm^2

OpenStudy (anonymous):

omg ur a life savior right now

OpenStudy (anonymous):

i m kinda tired right now lol can you solve the third part ???

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

ok you just have to find the ratios of their area and volume

OpenStudy (anonymous):

Maybe you'd like a reason for the first answer.. http://www.andrews.edu/~calkins/math/webtexts/geom10.htm Lateral Area of a prism: perimeter × height Since the perimeter is a linear function of length + width, the perimeter is scaled by r, and the height is scaled by r as well, so there will be a factor of r^2 in the final answer. What is \[(\frac{2}{3})^2\]